# Dexterous Contact-Rich Manipulation via the Contact Trust Region

### v1

@@ II-B Equations of Motion
$\mathbf{P}(q){q_{\texttt{+}}}+b(q,u)-\sum_{i=1}^{n_{c}}\mathbf{J}(q)^{\top}\lambda_{i}=0,\;\text{where}$ (3a)
@@
II-C Contact Constraints
Additional constraints further qualify the relationship between the next configurations ${q_{\texttt{+}}}$ and contact forces $\lambda_{i}$:
  1. 1.
    non-penetration: next configuration is non-penetrating,
  2. 2.
    friction cone: $\lambda_{i}$ must be inside the friction cone,
  3. 3.
    complementarity: contact cannot be applied from a distance, and the direction of friction opposes the movement (i.e. maximum dissipation).
@@ II-C Contact Constraints
Anistescu’s relaxation [51] allows us to write down these contact constraints in a convex manner by introducing a mild non-physical artifact, where sliding in the tangential plane results in separation in the normal direction,
@@
II-D Contact Dynamics as Convex Optimization
We can frame the equations of motion with contact as a search for the next configurations ${q_{\texttt{+}}}$ that satisfy the constraints introduced in Section II-B and Section II-C,
@@ II-D Contact Dynamics as Convex Optimization
where $\mathcal{I}_{c}$ is the index set over potential contact pairs. Note that stationarity (3), primal feasibility (5a), dual feasibility (5b), and complementary slackness (5c) are the KKT conditions of this SOCP. For complementary slackness of SOCPs, we note that we do not impose element-wise slackness: the primal and dual vectors can both be non-zero as long as they are orthogonal [54, 11].
@@
II-E Smoothing of Contact Dynamics
Despite being continuous and piecewise smooth, the contact dynamics map (8) is well known to have discontinuous gradients [7], complicating the application of planning and control methods that rely on local smooth approximations. More specifically, gradient-based methods can be destabilized by fast-changing gradients induced by quickly switching between smooth pieces of the contact dynamics (also called contact modes) [18]. In addition, the gradient landscape can have flat regions if none of the contact modes are active, leaving optimizers stranded without informative directions of improvement [9]. As such, various dynamic smoothing methods have been proposed to relax this numerical difficulty [12, 11, 18, 7, 26, 55].
In [12, 11], a method of smoothing out optimization-based dynamics was introduced based on the log-barrier (interior-point) relaxation of (7), which solves for the KKT equations where the complementarity condition is perturbed by a positive constant $\kappa$  [^1] The constant 2 on the RHS of (9) is due to the squares inside the $\log$ in (10). For more details, see the degree of generalized logarithms [54, §11.6]. ,
@@ II-E Smoothing of Contact Dynamics
As the optimization (10) is an unconstrained convex program, its optimality conditions can be obtained by simply setting the gradient of the cost function w.r.t. $q_{+}$ to 00:
@@
As an unconstrained optimization problem, (10) technically does not have dual variables. However, if we consider solving the SOCP dynamics (7) with the interior point method, (10) corresponds to one major iteration along the central path [54, §11.6]. Therefore, we can identify average contact forces with the dual feasible points of (7) along the central path:
@@
II-F Sensitivity Analysis of Smoothed Contact Dynamics
@@
The gradients can be computed through sensitivity analysis, which states that as we locally perturb the problem parameters, the solution of (14) change in a way that preserves (up to first-order) the optimality conditions of (14). This is equivalent to stating that the derivative of the equality constraints with respect to $q$ and $u$ are also zero at optimality. For instance, rewriting the derivative of (13) and (9) in matrix form gives
@@
where $\mathcal{H}_{i},\mathcal{V}_{i},\mathcal{D}_{i}$ stands for horizontal, vertical, and diagonal stacking of the terms from all contact pairs. Then, we can use the implicit function theorem to solve the system of equations and obtain the derivatives $\partial q_{+}/\partial u$ and $\partial\lambda_{i}/\partial u$ (more details in Section X-A). The derivatives with respect to $q$ may be obtained in a similar fashion, but are more involved due to the $\partial\mathbf{J}_{i}/\partial q$ term (curvature).
III Local Approximation of Contact Dynamics
@@
A majority of first-order algorithms for optimization and optimal control rely on iteratively making a local approximation of the problem that is more amenable for computation. For instance, gradient-descent makes a linear approximation, and Newton’s method makes a quadratic one. However, we rarely handle the case where the constraint itself is the result of an optimization. In this section, we answer the question: what is the right local approximation of the contact dynamics that i) is computationally convenient, and ii) captures the correct local behavior?
III-A Linear Model over Smoothed Primal and Dual Variables
@@
However, many previous works [9, 12, 11, 56, 7] have noted that due to the non-smooth nature of contact, it is beneficial to smooth the dynamics where a first-order Taylor approximation would be valid beyond the contact mode to which $(\bar{q},\bar{u})$ belongs [11]. A first-order Taylor expansion under such a smoothed dynamics model can be written as
@@
Then, this linear model satisfies the equality conditions of the perturbed SOCP (1), (9) to first order,
@@
with $\hat{\mathbf{P}}\coloneqq\mathbf{P}(\bar{q},\bar{u})+\frac{\partial\mathbf{P}}{\partial q}\delta q+\frac{\partial\mathbf{P}}{\partial u}\delta u$, and similarly for $\hat{b},\hat{\mathbf{J}},\hat{c}_{i}$.
@@
III-B Ellipsoidal Trust Region (ETR)
Lemma 1 implies that the linear model will no longer be accurate far away from $(\bar{q},\bar{u})$, even with the benefits of smoothing [11]. As such, many optimization methods are further concerned with where we can trust the local model, and introduce the concept of a trust region around ($\bar{q},\bar{u})$, which further qualifies the accuracy of the linear model [21, 23]. To simplify computation, these trust regions are often chosen to be simple geometric primitives.
@@ III-B Ellipsoidal Trust Region (ETR)
III-C The Contact Trust Region (CTR)
However, contact is more special in that it is a unilaterally constrained dynamical system [^2] Bilateral constraints are often included in sensitivity analysis, so would be implicitly obeyed by the linear model.. If points in the trust region predicts a behavior of primal and dual variables that are not feasible, this can also cause a large discrepancy in the approximated model within the trust region. In practice solvers would also try to find an ellipsoidal trust region such that all the elements within the trust region would result in a feasible prediction [57, 58]. However, this is often achieved by shrinking the ETR centered at the linearization point, resulting in overly conservative(small) trust regions.
@@ III-C The Contact Trust Region (CTR)
[…]
@@
III-D The Relaxed Contact Trust Region (R-CTR)
@@
[…]
@@
In both Definition 3 and Definition 4, we keep $\hat{q}^{\mathrm{a}}$ although they are not directly used in the trust region definitions: they are useful for imposing joint and torque limits, which are detailed in Section III-F.
@@
$\mathcal{M}_{\mathbf{\Sigma,\kappa}}(\bar{q},\bar{u})\coloneqq\{\hat{q}_{+}|$ $\hat{q}_{+}=\mathbf{A}_{\kappa}\delta q+\mathbf{B}_{\kappa}\delta u+f_{\kappa}(\bar{q},\bar{u}),$ (28a)
@@
$\mathcal{M}^{\mathcal{A}}_{\mathbf{\Sigma,\kappa}}(\bar{q},\bar{u})\coloneqq\{\hat{q}_{+}|$ $\hat{q}_{+}=\mathbf{B}_{\kappa}\delta u+f_{\kappa}(\bar{q},\bar{u}),$ (29a)
@@
III-E Mechanics Derivation of the Motion Set
We can further motivate the idea that relaxing the primal feasibility constraints improves local approximations of feasible object motions by connecting the RA-CTR to classical constructs in manipulation, such as the motion cone [59, 60] and the wrench set [61, 62]. In this section, we establish these connections by providing an alternative derivation of the action-only motion set (Definition 6), starting from contact forces instead of Taylor expansions.
@@ III-E Mechanics Derivation of the Motion Set
The Generalized Friction Cone. The generalized friction cone [63] is a linear mapping of the friction cone to the space of object wrenches. This can be obtained by applying the contact Jacobian that linearizes the kinematics of the contact point with respect to object coordinates,
@@
The Wrench Set. The wrench set [61, 62] is defined as the set of all achievable wrenches that can be applied to an object given all possible contact forces that can be applied from a given configuration. This quantity has classically served as an important metric in grasping analysis [61, 5, 64]. While the classical wrench set has bounds on the contact force, we instead apply bounds on the actuator input ($\|\delta u\|_{\mathbf{\Sigma}}^{2}\leq 1$).
@@
The Motion Set. The motion set is analogous to motion cones [59, 60]. While classical motion cones describe the set of feasible object velocity under a single patch contact, we define motion cone as the set of all feasible displacements achievable by multiple contacts in one time step.
@@ III-F1 Joint Limits
III-F2 Torque Limits
We often want to limit the torque applied by the robot to the object by imposing lower and upper bounds $\tau_{lb}$ and $\tau_{ub}$ on joint torque. Although the quasi-dynamic formulation cannot account for dynamic transient torques, we can compute the steady-state torque experienced by the robot using the difference in the sent position command and the predicted position [65]. Thus, the set of achievable $\delta u$ that does not violate this steady-state torque limit can be written as
@@ III-F2 Torque Limits
IV Local Planning and Control
We now study local gradient-based planning and control, which is one of the core applications of our proposed contact trust region. Among possible formulations of this problem, we focus on coming up with a configuration and input sequence that moves the system towards a goal configuration $q_{\text{goal}}$. The full nonlinear form of this problem can be written as
@@ IV Local Planning and Control
One of the biggest challenges in solving (36) is handling the non-smooth contact dynamics constraint (36b). Although MIP-based methods [28, 29] have struggled to scale up to complex, contact-rich problems, the MIP formulation reveals why the problem is hard: the search through the exponentially many contact modes.
In this section, we present a method for solving (36) by incorporating contact dynamics smoothing and the contact trust region into an iLQR-like[39] trajectory optimization scheme. Through two toy problems, we show that the proposed method can iteratively approach an advantageous contact mode for reaching the given goals, even when the initial guess is in a contact mode with non-informative gradient. In addition, we present how the method can be used as a model predictive controller and be extended to control contact forces.
IV-A Trajectory Optimization with R-CTR
@@
[…]
@@
Here (37c) is the standard linear dynamics constraint in linear MPC, where $\delta q_{t+1}\coloneqq q_{t+1}-f(\bar{q}_{t},\bar{u}_{t})$; $\tilde{\mathcal{S}}_{\mathbf{\Sigma,\kappa}}$ in (37d) is the R-CTR (26); (37c) and (37d) constitute our local approximation of the contact dynamics constraint (36b); (37h) is due to the initial condition constraint (36d). The sub trajectory optimization problem (37) is a standard SOCP that can be solved by off-the-shelf conic solvers. We note that other constraints as joint limits and torque limits in Section III-F can be added trivially to this formulation by incorporating them into $\tilde{\mathcal{S}}_{\mathbf{\Sigma},\kappa}$.
@@
Lastly, we impose the relaxed CTR constraint in (37d) instead of the full CTR, as the full CTR often leads to under approximation of the 1-step reachable set of the true dynamics. We have shown this via simple examples in Section III, and will show more experimental support for the choice of using R-CTR in Section V.

### v3

@@ 2.2 Equations of Motion
$\mathbf{P}(q){q_{\texttt{+}}}+b(q,u)-\sum_{i=1}^{n_{c}}\mathbf{J}_{i}(q)^{\top}\lambda_{i}=0,\;\text{where}$ (3a)
@@
2.3 Contact Constraints
Additional constraints further qualify the relationship between the next configurations $q_{\texttt{+}}$} and contact forces $\lambda_{i}$: i) non-penetration: next configuration is non-penetrating; ii) friction cone: $\lambda_{i}$ must be inside the friction cone; iii) complementarity: contact cannot be applied from a distance, and the direction of friction opposes the movement (i.e. maximum dissipation).
@@ 2.3 Contact Constraints
Anistescu’s relaxation (Anitescu, 2006) allows us to write down these contact constraints in a convex manner by introducing a mild non-physical artifact, where sliding in the tangential plane results in separation in the normal direction,
@@
2.4 Contact Dynamics as Convex Optimization
We can frame the equations of motion with contact as a search for the next configurations $q_{\texttt{+}}$} that satisfy the constraints introduced in Section 2.2 and Section 2.3,
@@ 2.4 Contact Dynamics as Convex Optimization
where $\mathcal{I}_{c}$ is the index set over potential contact pairs. Note that stationarity (3), primal feasibility (5a), dual feasibility (5b), and complementary slackness (5c) are the KKT conditions of this SOCP. For complementary slackness of SOCPs, we note that we do not impose element-wise slackness: the primal and dual vectors can both be non-zero as long as they are orthogonal (Boyd and Vandenberghe, 2004; Pang et al., 2023).
@@
2.5 Smoothing of Contact Dynamics
Despite being continuous and piecewise smooth, the contact dynamics map (8) is well known to have discontinuous gradients (Posa et al., 2014), complicating the application of planning and control methods that rely on local smooth approximations. More specifically, gradient-based methods can be destabilized by fast-changing gradients induced by quickly switching between smooth pieces of the contact dynamics (also called contact modes) (Suh et al., 2022b). In addition, the gradient landscape can have flat regions if none of the contact modes are active, leaving optimizers stranded without informative directions of improvement (Suh et al., 2022a). As such, various dynamic smoothing methods have been proposed to relax this numerical difficulty (Howell et al., 2023a; Pang et al., 2023; Suh et al., 2022b; Posa et al., 2014; Onol et al., 2019; Le Lidec et al., 2024).
In Howell et al. (2023a); Pang et al. (2023), a method of smoothing out optimization-based dynamics was introduced based on the log-barrier (interior-point) relaxation of (7), which solves for the KKT equations where the complementarity condition is perturbed by a positive constant $\kappa$  [^1] The constant 2 on the RHS of (9) is due to the squares inside the $\log$ in (10). For more details, see the degree of generalized logarithms (Boyd and Vandenberghe, 2004, §11.6). ,
@@ 2.5 Smoothing of Contact Dynamics
As the optimization (10) is an unconstrained convex program, its optimality conditions can be obtained by simply setting the gradient of the cost function w.r.t. $q_{+}$ to 0:
@@
As an unconstrained optimization problem, (10) technically does not have dual variables. However, if we consider solving the SOCP dynamics (7) with the interior point method, (10) corresponds to one major iteration along the central path (Boyd and Vandenberghe, 2004, §11.6). Therefore, we can identify average contact forces with the dual feasible points of (7) along the central path:
@@
2.6 Sensitivity Analysis of Smoothed Contact Dynamics
@@
The gradients can be computed through sensitivity analysis, which states that as we locally perturb the problem parameters, the solution of (14) change in a way that preserves (up to first-order) the optimality conditions of (14). This is equivalent to stating that the derivative of the equality constraints with respect to $q$ and $u$ are also zero at optimality. For instance, rewriting the derivative of (13) and (9) in matrix form ($\nu_{i}$ and $c_{i}$ need to be expanded using their definitions in (5a) and (7c), respectively) gives
@@
where $\mathcal{H}_{i},\mathcal{V}_{i},\mathcal{D}_{i}$ stands for horizontal, vertical, and diagonal stacking of the terms from all contact pairs. Then, we can use the implicit function theorem to solve the system of equations and obtain the derivatives $\partial q_{+}/\partial u$ and $\partial\lambda_{i}/\partial u$ (more details in Section A.1). The derivatives with respect to $q$ may be obtained in a similar fashion, but are more involved due to the $\partial\mathbf{J}_{i}/\partial q$ term (curvature).
3 Local Approximation of Contact Dynamics
@@
A majority of first-order algorithms for optimization and optimal control rely on iteratively making a local approximation of the problem that is more amenable for computation. For instance, gradient-descent makes a linear approximation, and Newton’s method makes a quadratic one. However, we rarely handle the case where the constraint itself is the result of an optimization. In this section, we answer the question: what is the right local approximation of the contact dynamics that (i) is computationally convenient, and (ii) captures the correct local behavior?
3.1 Linear Model over Smoothed Primal and Dual Variables
@@
However, many previous works (Suh et al., 2022a; Howell et al., 2023a; Pang et al., 2023; Shirai et al., 2024b; Posa et al., 2014) have noted that due to the non-smooth nature of contact, it is beneficial to smooth the dynamics where a first-order Taylor approximation would be valid beyond the contact mode to which $(\bar{q},\bar{u})$ belongs (Pang et al., 2023). A first-order Taylor expansion under such a smoothed dynamics model can be written as
@@
Then, this linear model satisfies (13) and (9), the optimality conditions of the perturbed SOCP (10), to first order:
@@
with $\hat{\mathbf{P}}\coloneqq\mathbf{P}(\bar{q},\bar{u})+\frac{\partial\mathbf{P}}{\partial q}\delta q+\frac{\partial\mathbf{P}}{\partial u}\delta u$, and similarly for $\hat{b}$, $\hat{\mathbf{J}}$, and $\hat{c}_{i}$.
@@
3.2 Ellipsoidal Trust Region (ETR)
Lemma 1 implies that the linear model will no longer be accurate far away from $(\bar{q},\bar{u})$, even with the benefits of smoothing (Pang et al., 2023). As such, many optimization methods are further concerned with where we can trust the local model, and introduce the concept of a trust region around ($\bar{q},\bar{u})$, which further qualifies the accuracy of the linear model (Sorensen, 1982; More, 1993). To simplify computation, these trust regions are often chosen to be simple geometric primitives.
@@ 3.2 Ellipsoidal Trust Region (ETR)
3.3 Contact Trust Region (CTR)
However, contact is more special in that it is a unilaterally constrained dynamical system [^2] Bilateral constraints are often included in sensitivity analysis, so would be implicitly obeyed by the linear model.. If points in the trust region predicts a behavior of primal and dual variables that are not feasible, this can also cause a large discrepancy in the approximated model within the trust region. In practice solvers would also try to find an ellipsoidal trust region such that all the elements within the trust region would result in a feasible prediction (Moré and Sorensen, 1983; Niu and Yuan, 2010). However, this is often achieved by shrinking the ETR centered at the linearization point, resulting in overly conservative(small) trust regions.
@@ 3.3 Contact Trust Region (CTR)
$\mathcal{S}_{\mathbf{\Sigma},\kappa}(\bar{q},\bar{u})\coloneqq\{$ $(\delta q,\delta u)|\delta z^{\top}\mathbf{\Sigma}\delta z\leq 1,\delta z=(\delta q,\delta u),$ (24a)
@@
Notably, the convex constraints (24b)-(24e) are locally equivalent to the perturbed (smoothed) KKT conditions of contact dynamics (6) (which is equivalent to (7)). By Lemma 1, the primal and dual Taylor expansions, (24b) and (24c), are equivalent to the stationarity condition (13) (which is equivalent to (3)) and the perturbed complementarity condition (9) (which is the perturbation of (5c)). Moreover, primal feasibility (24d) locally implies (5a), and dual feasibility (24e) implies (5b).
@@
3.4 Relaxed Contact Trust Region (R-CTR)
@@
$\tilde{\mathcal{S}}_{\mathbf{\Sigma},\kappa}(\bar{q},\bar{u})\coloneqq\{$ $(\delta q,\delta u)|\delta z^{\top}\mathbf{\Sigma}\delta z\leq 1,\delta z=(\delta q,\delta u),$ (26a)
@@
In both Definition 3 and Definition 4, we keep $\hat{q}^{\mathrm{a}}$ although they are not directly used in the trust region definitions: they are useful for imposing joint and torque limits, which are detailed in Section 3.6.
@@
[…]
@@
[…]
@@
3.5 Mechanics Derivation of the Motion Set
We can further motivate the idea that relaxing the primal feasibility constraints improves local approximations of feasible object motions by connecting the RA-CTR to classical constructs in manipulation, such as the motion cone (Mason, 1986; Chavan-Dafle et al., 2020) and the wrench set (Ferrari and Canny, 1992; Lynch and Park, 2017). In this section, we establish these connections by providing an alternative derivation of the action-only motion set (Definition 6), starting from contact forces instead of Taylor expansions.
@@ 3.5 Mechanics Derivation of the Motion Set
The Generalized Friction Cone. The generalized friction cone (Erdmann, 1994) is a linear mapping of the friction cone to the space of object wrenches. This can be obtained by applying the contact Jacobian that linearizes the kinematics of the contact point with respect to object coordinates,
@@
The Wrench Set. The wrench set (Ferrari and Canny, 1992; Lynch and Park, 2017) is defined as the set of all achievable wrenches that can be applied to an object given all possible contact forces that can be applied from a given configuration. This quantity has classically served as an important metric in grasping analysis (Ferrari and Canny, 1992; Han et al., 2000; Dai et al., 2016). While the classical wrench set has bounds on the contact force, we instead apply bounds on the actuator input ($\|\delta u\|_{\mathbf{\Sigma}}^{2}\leq 1$).
@@
The Motion Set. The motion set is analogous to motion cones (Mason, 1986; Chavan-Dafle et al., 2020). While classical motion cones describe the set of feasible object velocity under a single patch contact, we define motion cone as the set of all feasible displacements achievable by multiple contacts in one time step.
@@ 3.6.1 Joint Limits
3.6.2 Torque Limits
We often want to limit the torque applied by the robot to the object by imposing lower and upper bounds $\tau_{lb}$ and $\tau_{ub}$ on joint torque. Although the quasi-dynamic formulation cannot account for dynamic transient torques, we can compute the steady-state torque experienced by the robot using the difference in the sent position command and the predicted position (Pang and Tedrake, 2022). Thus, the set of achievable $\delta u$ that does not violate this steady-state torque limit can be written as
@@ 3.6.2 Torque Limits
4 Local Planning and Control
We now study local gradient-based planning and control, which is one of the core applications of our proposed contact trust region. Among possible formulations of this problem, we focus on computing a configuration and input sequence that moves the system towards a goal configuration $q_{\text{goal}}$. The full nonlinear form of this problem can be written as
@@ 4 Local Planning and Control
One of the biggest challenges in solving (36) is handling the non-smooth contact dynamics constraint (36b). Although MIP-based methods (Marcucci et al., 2017; Marcucci and Tedrake, 2019) have struggled to scale up to complex, contact-rich problems, the MIP formulation reveals why the problem is hard: the search through the exponentially many contact modes.
In this section, we present a method for solving (36) by incorporating contact dynamics smoothing and the contact trust region into an iLQR-like (Li and Todorov, 2004) trajectory optimization scheme. Through two toy problems, we show that the proposed method can iteratively approach an advantageous contact mode for reaching the given goals, even when the initial guess is in a contact mode with non-informative gradient. In addition, we present how the method can be used as a model predictive controller.
4.1 Trajectory Optimization with R-CTR
@@
$\text{s.t.}\quad\delta q_{t+1}=\mathbf{A}_{\kappa,t}\delta q_{t}+\mathbf{B}_{\kappa,t}\delta u_{t},\;t=0\dots T-1,$ (37c)
@@
Here (37c) is the standard linear dynamics constraint in linear MPC, where $\delta q_{t+1}\coloneqq q_{t+1}-f(\bar{q}_{t},\bar{u}_{t})$; $\tilde{\mathcal{S}}_{\mathbf{\Sigma,\kappa}}$ in (37d) is the R-CTR (26); (37c) and (37d) constitute our local approximation of the contact dynamics constraint (36b); (37h) is due to the initial condition constraint (36d). The sub trajectory optimization problem (37) is a standard SOCP that can be solved by off-the-shelf conic solvers. We note that other constraints as joint limits and torque limits in Section 3.6 can be added trivially to this formulation by incorporating them into $\tilde{\mathcal{S}}_{\mathbf{\Sigma},\kappa}$.
@@
Lastly, we impose the relaxed CTR constraint in (37d) instead of the full CTR, as the full CTR often leads to under approximation of the 1-step reachable set of the true dynamics. We have shown this via simple examples in Section 3, and will show more experimental support for the choice of using R-CTR in Section 5.